Shift report 20000814 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Aug 14 2000 - 16:04:41 EDT

  • Next message: Peter H. L. Christiansen: "Revised tpc monitor"

    ------------------------------------------------------------
    Date: Mon Aug 14 16:04:41 2000, Shift: 08:00-16:00
    Supervisor: hb, with: ro
    Log book#4 pages 87 - 90
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was available from 12:15 to 13:45. Beam 
    intensities were 7087 in Blue and 1719 in Yellow.
    
    2. Run 2251 was startet and finnished. MRS at 90 deg. 
    and 4KGauss (700A) fields. FFS is the same as before.
    89907 collisions were contained.
    
    3. T2 readout was broken. RD fixed it. Turned it on
    12:35
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Aug 14 2000 - 16:08:10 EDT